# Nonlinear Operator Learning Using Energy Minimization and MLPs

### v1

@@ 4.3 Nonlinear elasticity
A promising application of solution operator networks is to use the output as an initial guess for Newton’s method when solving nonlinear problems, see, e.g., [5, 10, 2]. This is a complementary way of using machine learning for PDE solving rather than as a complete solver method, i.e., to consider the operator network’s solution the finished product. Fair criticism towards the latter includes potential loss of accuracy when compared to well-established methods such as FEM. Using the network solution as initial guess for Newton’s method is a way to preserve accuracy and at the same time speed up the iterative process.
@@
Here we consider a nonlinear neo-Hookean elasticity problem: a two dimensional cantilever beam under the influence of a given external force. The external force is defined by two variables: the midpoint of a fixed-length surface interval of application on top of the beam and the vertical magnitude. The solution to the problem is the displacement field over the beam. We generate a structured uniform triangular 40x2 mesh of the resting beam and consider standard P1 finite elements. This results in 240 DOFs, since every mesh node ($41\cdot 3=123$) has two DOFs (one for each dimension) and the DOFs of the left-most three nodes are removed since that is where the beam is fixed.
@@
The solution operator we aim to learn is thus the map from the 2 force variables to the 240 finite element DOFs giving the corresponding displacement of the beam. We use a simple MLP architecture that takes 2 input variables (the force variables), has 4 hidden layers of width 256 with ELU activation and one output layer of width 240 (number of DOFs) with no activation. Just as in the previous examples we use the energy functional corresponding to the PDE as the foundation for the loss function.
@@
During training, we pick 32 randomly selected tuples of force variables in each iteration to compute an average loss. The midpoint of the interval of application is sampled uniformly along the beam and the magnitude is sampled from a standard normal distribution. The optimization is performed with the Adam optimizer where we perform $10^{6}$ iterations with a decreasing learning rate. The learning rate starts at 1e-4 and after every 250k iterations it is decreased by a factor of 0.5.
@@
We consider the specific problem of computing the displacement when the external force is located at the right end of the beam and directed downwards with magnitude 1.0. We perform the Newton iterations with a FEniCS implementation of the problem. All computations are performed on an Apple M1 CPU. Results from using the network solution and a standard initial guess (the zero function) as input for Newton’s method are presented in Table 8 and Figure 3.
@@
Table 8: Computational times for using different initial guesses for Newton’s method.
Initial guess $E$(solution) Inference time Newton solver time Total time
Network solution -0.014397 9.4e-5 s 0.015678 s 0.015772 s
@@
(a) Network solution as initial guess.
@@
Figure 3: Initial guesses for Newton’s method with resulting converged solutions. The converged solution is the same in both cases.
@@
The results in Table 8 and Figure 3 show that using the network solution as initial guess for Newton’s method preserves accuracy while at the same time provides a speed-up compared to using the standard zero function as initial guess. We note that we have not taken the training time of the network into account, so to properly benefit from using the network solution as initial guess a large enough number of problems will have to be solved so that all individual problem gains compensate for the training time.

### v2

@@ 4.3 Nonlinear elasticity
A promising application of solution operator networks is to use the output as an initial guess for Newton’s method when solving nonlinear problems, see, e.g., [16, 28, 2]. This is a complementary way of using machine learning for PDE solving rather than as a complete solver method, i.e., to consider the operator network’s prediction the finished product. Fair criticism towards the latter includes potential loss of accuracy when compared to standard methods such as FEM. Using the network prediction as initial guess for Newton’s method is a way to preserve accuracy while at the same time potentially speed up the iterative process. It is also a natural example of when machine learning frameworks can be combined with existing software for standard methods, thus motivating the learning of the standard method’s approximation, here the final element solution. The finite element software we use here is the popular finite element library FEniCS [23].
@@
Using neural networks to predict hyperelasticity solutions has been done in, e.g., [27] with a discretization-free method and in [7] with a meshfree collocation method. Here we consider a nonlinear neo-Hookean elasticity problem: a two dimensional cantilever beam under the influence of a given external force. The external force is defined by three variables: the midpoint $p$ of a fixed-length surface interval of application on top of the beam and the horizontal and vertical magnitudes $F_{x}$ and $F_{y}$, respectively. The solution to the problem is the displacement field over the beam. We generate a structured uniform triangular 40x2 mesh of the resting beam and consider standard P1 finite elements. This results in 240 DOFs, since every mesh node ($41\cdot 3=123$) has two DOFs (one for each dimension) and the DOFs of the left-most three nodes are removed since that is where the beam is fixed.
@@
The solution operator we aim to learn is thus the map from the 3 force variables to the 240 finite element DOFs giving the corresponding displacement of the beam. We use a simple MLP architecture that takes 3 input variables (the force variables), has 4 hidden layers of width 256 with ELU activation and one output layer of width 240 (number of DOFs) with no activation. Just as in the previous examples we use the energy functional corresponding to the PDE as the foundation for the loss function. Here, the energy is
$E(v)=\int_{\Omega}\Psi(v)\,\text{dx}-\int_{\partial\Omega\cap\text{supp}(F)_{p}}F\cdot v\,\text{ds}$ (4.21)
where $\Psi$ is the neo-Hookean elastic stored energy density and $F=(F_{x},F_{y})$ is the external force field. Gravity is not included.
@@
During training, we pick 32 randomly selected tuples of force variables in each iteration to compute an average loss. The midpoint of the interval of application is sampled uniformly along the beam $p\sim\mathcal{U}=\mathcal{U}([x_{\min},x_{\max}])$, where $x_{\min}$ and $x_{\max}$ are the left and right endpoints of the beam, respectively. The horizontal and vertical magnitudes are sampled from normal distributions with mean 0 and standard deviation $\sigma$, i.e., $F_{x},F_{y}\sim\mathcal{N}(0,\sigma^{2})$. The optimization is performed with the Adam optimizer where we perform $10^{6}$ iterations with a decreasing learning rate. The learning rate starts at 1e-4 and after every 250k iterations it is decreased by a factor of 0.5.
Sampling the force magnitudes from normal distributions with mean 0 means that we train the networks more for problems with solutions close to the zero function, i.e., the unbent beam. It is thus reasonable to expect the networks to perform better for such problems than for problems with large forces resulting in large deformations. The previously used relative learning errors are thus unsuitable here, since they could potentially give very large errors for problems with solutions close to zero and smaller errors for more deformed solutions. Letting $\boldsymbol{F}=(p,F_{x},F_{y})$ and $|\Omega|$ denote the Lebesgue measure of $\Omega$, we therefore instead compute the $\Omega$-relative learning errors
$|E(\mathcal{A}_{h,\theta}(\boldsymbol{F}_{i}))-E(\mathcal{A}_{h}(\boldsymbol{F}_{i}))|/|\Omega|$ (4.22)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{F}_{i})-\mathcal{A}_{h}(\boldsymbol{F}_{i})\|_{L^{2}}/|\Omega|^{1/2}$ (4.23)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{F}_{i})-\mathcal{A}_{h}(\boldsymbol{F}_{i})\|_{H_{0}^{1}}/|\Omega|^{1/2}$ (4.24)
We consider two different force situations: basic beam bending and extreme bending.
4.3.1 Basic bending
Here, we only consider one simple sampling case for the force variables in which we do not include a horizontal magnitude. The force variables are sampled by
$p\sim\mathcal{U}\quad F_{x}=0\quad F_{y}\sim\mathcal{N}(0,1)$ (4.25)
Training and inference times for the basic bending case on both an A100 GPU and an Apple M1 CPU are presented in Table 9. The training times clearly show the benefit of GPU training.
Table 9: Training and inference times for the basic bending case.
Hardware Training time Inference time
GPU (A100) 558 s 0.8 ms
CPU (M1) 2393 s 0.13 ms
Learning errors for the basic bending case are presented in Table 10 and corresponding histogram plots are presented in Figure 4. The histograms show that the error distributions are skewed towards zero, indicating some degree of well-trainedness.
Table 10: $\Omega$-relative learning errors in different norms for the basic bending case. Sample size $K=10^{3}$.
Training & test data Measure Energy $E(\cdot)$ $L^{2}$-norm $H_{0}^{1}$-norm
$p\sim\mathcal{U}\quad F_{x}=0\quad F_{y}\sim\mathcal{N}(0,1)$ Mean 0.0058 0.0115 0.0477
Std 0.0085 0.0126 0.0481
(a) Energy
(b) $L^{2}$-norm
(c) $H_{0}^{1}$-norm
Figure 4: Histogram plots for the different $\Omega$-relative learning errors for the basic bending case, i.e., Table 10. Sample size $K=10^{3}$.
@@ 4.3.1 Basic bending
We now consider the specific problem of computing the displacement when the external force is located at the right end of the beam and directed downwards with magnitude 1.0. We use the operator network for the basic bending case and perform the Newton iterations with a FEniCS implementation of the problem. All computations are performed on an Apple M1 CPU. Results from using the network prediction and a standard initial guess (the zero function) as input for Newton’s method are presented in Table 11 and Figure 5.
@@
Table 11: Computational times on an Apple M1 CPU for using different initial guesses for Newton’s method.
Initial guess $E(u_{h})$ Inference time Newton solver time Total time
Network prediction -0.014397 9.4e-5 s 0.015678 s 0.015772 s
@@
(a) Network prediction as initial guess.
@@
Figure 5: Different initial guesses for Newton’s method with resulting converged solutions for a basic bending case. The converged solution $u_{h}$ is the same in both cases.
@@
The results in Table 11 and Figure 5 show that using the network prediction as initial guess for Newton’s method preserves accuracy while at the same time provides a speed-up compared to using the standard zero function as initial guess. However, to properly benefit from using the network prediction as initial guess, a large enough number of problems will have to be solved so that all individual problem gains compensate for the training time. From Table 9, we compute that one needs to solve at least 81,034 problems with GPU training and 347,517 with CPU training to get this benefit. This might sound like a lot, but if the trained network is incorporated into software intended for many users or if many instances have to be solved, e.g., in an optimization process or for a time-dependent version of the problem, the numbers quickly diminish in perspective.
4.3.2 Extreme bending
Here, we consider three more advanced sampling cases for the force variables:
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,1)$ (4.26)
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,4)$ (4.27)
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,16)$ (4.28)
Training and inference times for the extreme bending cases on both an A100 GPU and an Apple M1 CPU are presented in Table 12. Again the results show the benefits of GPU training.
Table 12: Training and inference times for the extreme bending cases. The A100 GPU Utilization was 46% in all three cases.
Training & test data Hardware Training time Inference time
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,1)$ GPU (A100) 583 s 0.85 ms
CPU (M1) 2398 s 0.17 ms
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,4)$ GPU (A100) 583 s 0.88 ms
CPU (M1) 2438 s 0.14 ms
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,16)$ GPU (A100) 582 s 0.86 ms
CPU (M1) 2672 s 0.13 ms
Learning errors for the extreme bending cases are presented in Table 13 and histograms for the last case are presented in Figure 6. The learning errors increase with increased variance of the data distributions. This is reasonable since wider distributions can be thought of as larger domains of the map that the network approximates. Larger domains, mean more complicated maps and in turn harder to train networks that are harder to train well. Nevertheless, the histograms for the last case all show small error distributions skewed towards zero, indicating some degree of well-trainedness.
Table 13: $\Omega$-relative learning errors in different norms for the extreme bending cases. Sample size $K=10^{3}$.
Training & test data Measure Energy $E(\cdot)$ $L^{2}$-norm $H_{0}^{1}$-norm
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,1)$ Mean 0.0443 0.0395 0.1564
Std 0.1720 0.0749 0.2928
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,4)$ Mean 0.1396 0.0574 0.2236
Std 0.4489 0.1000 0.3819
$p\sim\mathcal{U}\quad F_{x},F_{y}\sim\mathcal{N}(0,16)$ Mean 0.2299 0.0704 0.2669
Std 0.7610 0.1104 0.4052
(a) Energy
(b) $L^{2}$-norm
(c) $H_{0}^{1}$-norm
Figure 6: Histogram plots for the different $\Omega$-relative learning errors for the last extreme bending case, i.e., the last row of Table 13. Sample size $K=10^{3}$.
We now consider the specific problem of computing the displacement when the external force is again located at the right end of the beam but successively turns around the fixed end, forcing the beam to curl. This is different from the basic bending example since the force now changes when the beam has been bent to a certain degree. Doing this with the standard approach means performing sets of Newton iterations: From an initial state of the beam and an initial force, Newton iterations are performed until convergence. The converged state is then used as the new starting state together with a new force, and so on. The last converged state is the desired one and the other ones are simply intermediates. The forces we consider for the standard approach are:
$F_{1}=(0.0,-0.99)\quad F_{2}=(-1.5,0.0)\quad F_{3}=(0.0,5.0)$ (4.29)
This means that we will get three states: two intermediates and one final curled state. For the machine learning approach, we note that the operator networks we have trained cannot reach the final curled state of the beam. This is because the forces used as input to the network are all applied to the unbent beam. However, the operator networks can still provide predictions of intermediate states that can be useful as initial guesses. We use an operator network from the last extreme bending case, i.e., where $F_{x},F_{y}\sim\mathcal{N}(0,16)$ during training, and take the network intermediate prediction coming from $F=(-10.0,-5.0)$ as initial guess. We again perform the Newton iterations with a FEniCS implementation of the problem. All computations are performed on an Apple M1 CPU. Results from using the network intermediate state and the successive beam configurations starting from the standard initial guess (the zero function) as input for Newton’s method are presented in Table 14 and Figure 7.
Table 14: Computational times on an Apple M1 CPU for using different initial guesses for Newton’s method. The network intermediate state comes from an operator network where $F_{x},F_{y}\sim\mathcal{N}(0,16)$ during training.
Initial guess $E(u_{h})$ Inference time Newton solver time Total time
Network intermediate 0.067984 9.5e-5 s 0.019925 s 0.02002 s
Zero function 0.067984 0.04784 s 0.04784 s
(a) Network intermediate as initial guess. The unbent beam is present for reference.
(b) Zero function (unbent beam) as initial guess. The two intermediate states are also present.
Figure 7: Different initial guesses for Newton’s method with resulting converged solutions for an extreme bending case. The converged solution $u_{h}$ is the same in both cases. The network intermediate state comes from an operator network where $F_{x},F_{y}\sim\mathcal{N}(0,16)$ during training.
The results in Table 14 and Figure 7 show that just as in the basic bending case, using the network output as initial guess for Newton’s method preserves accuracy while at the same time provides a speed-up compared to using the standard zero function as initial guess. Furthermore, the gain from using the network output is higher in the extreme bending case compared to the basic one. This is easily seen by comparing the total times in Table 11 and 14. For the basic bending case we have that the total time with the network output is about 70% of the total time with the zero function. The corresponding quantity for the extreme bending case is 42%. This higher gain naturally affects the number of problems that have to be solved in order to compensate for the training time of the network. From the last case in Table 12, we compute that one needs to solve at least 20,921 problems with GPU training and 96,047 with CPU training to get a benefit from using the network output as initial guess when taking the training time into account. These numbers for the extreme bending case are 26% and 28% of those of the basic bending case for GPU and CPU training, respectively. We point out that the higher gain from using a network in an extreme bending case could in fact be even higher. This is so since here we only use the network to get an intermediate state, not a prediction of the solution as in the basic bending case. From Table 11 and 14, the Newton solver time for using a prediction in the basic bending case is 79% of the one for using an intermediate state in the extreme case, suggesting the even higher computational gain. As a final remark we point out that the two Newton solver times do not seem to differ that much while the corresponding beam configuration pairs (starting and converged states) differ quite a lot visually: From Figure 5 and 7, the starting state is much closer to the converged one in the basic bending case compared to in the extreme one. We take this to mean that most of the Newton computational work is performed towards the end, fine-tuning the beam configuration to the final state.
To summarize, combining operator networks with Newton’s method preserves accuracy and can provide a speed-up. In the case of nonlinear elasticity, the greater the deformation, the greater the potential speed-up from using operator networks.
5 Conclusions and outlook
We have presented a machine learning framework for learning solutions to a class of PDE problems. A core idea of the framework is to learn the corresponding discrete solution of some standard numerical method instead of aiming for the exact solution. The reason being that the standard method could be used to aid and enhance the framework. This core idea can in general be applied to various machine learning frameworks and standard methods but here we have considered a simple MLP-architecture together with energy minimization for the framework and FEM as the standard numerical method. We have presented both theoretical results (approximation error estimate) and practical applications (Newton’s method) that demonstrate how the framework may be beneficially combined with FEM. We have also presented pure framework results that show strengths and limitations of it as well as potential applications. These results are the learning errors, the usage of batches of elements for the energy during training, and the computation of quantities of interest.
Concerning avenues for future work, besides looking into more advanced network architectures and training algorithms, the last elasticity example provides a natural starting point. We note that in this example (the extreme bending case), the neural network is limited by the fact that the external forces are always applied to the initial state (the unbent beam). This could be improved by introducing time dependency where the forces are allowed to change during the bending process. This would also most likely mean that the network needs to take the current position of the beam as an input parameter. This is an interesting path for future research.